OpenStudy (anonymous):

Find 3x - y - 3z if x = -2, y = 1, and z = -2. -13 -1 1 13

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

put these values in eqn.... 3(-2)-1-3(-2) =-6-1+6 =-1 as multiplication oftwo nwgativ num is positiv...

OpenStudy (anonymous):

as multiplication of two negativ num is positive..* .